.....The Fermi velocity is_________



a) 2.32 x 10^6 m/s

b) 1.57 x 10^6m/s

C) 4.57 x 10^6m/s

D) 3.37 x 10^6 m/s
1
Expert's answer
2021-09-24T09:26:25-0400

For Fermi energy of copper 7 eV and, the Fermi velocity is


"v_F=\sqrt{\frac{2E_F}{m}}=\sqrt{\frac{2(7·1.602·10^{-19})}{9.109·10^{-31}}}=1.57·10^6\text{ m/s}."